HOME ESR EXPERIMENTS LABS etc
NucAR lab laser_lab Fritz-Bosch-Lab X-Ray Lab target Beavertail
  Beavertail  Not logged in ELOG logo
Message ID: 45     Entry time: Sat Apr 5 17:49:06 2025
Author: CG, GL, ID 
Category: Detectors 
Subject: Outside detector moved in  
After some Sergey-magic the outside pocket moves again. We slowly moved the outside detector in, we are clipping the beam at -20mm. Moved the detector to -40mm and take some data as of 17:48 (run0036.lmd)

The leakage currents of the SiPad detector jump back and forth 2 microamps (btw 25 and 4.5 microamps), the DSSD seems stable, the crystal too

https://elog.gsi.de/esr/E0203_E0018/207
ELOG V3.1.5-fc6679b